### v2

# BOTS: A Unified Framework for Bayesian Online Task Selection in LLM Reinforcement Finetuning

Abstract
Reinforcement finetuning (RFT) is a key technique for aligning Large Language Models (LLMs) with human preferences and enhancing reasoning, yet its effectiveness is highly sensitive to which tasks are explored during training. Uniform task sampling is inefficient, wasting computation on tasks that are either trivial or unsolvable, while existing task selection methods often suffer from high rollout costs, poor adaptivity, or incomplete evidence. We introduce BOTS, a unified framework for Bayesian Online Task Selection in LLM reinforcement finetuning. Grounded in Bayesian inference, BOTS adaptively maintains posterior estimates of task difficulty as the model evolves. It jointly incorporates explicit evidence from direct evaluations of selected tasks and implicit evidence inferred from these evaluations for unselected tasks, with Thompson sampling ensuring a principled balance between exploration and exploitation. To make implicit evidence practical, we instantiate it with an ultra-light interpolation-based plug-in that estimates difficulties of unevaluated tasks without extra rollouts, adding negligible overhead. Empirically, across diverse domains and LLM scales, BOTS consistently improves data efficiency and performance over baselines and ablations, providing a practical and extensible solution for dynamic task selection in RFT.  [^1] Our code is released at https://github.com/modelscope/Trinity-RFT/tree/main/examples/bots
11footnotetext: Equal contribution22footnotetext: Corresponding Authors. Email to {shenqianli.sql, yaliang.li, bolin.ding}@alibaba-inc.com
1 Introduction
Reinforcement finetuning (RFT) has become a key technique for aligning Large Language Models (LLMs) with human preferences and enhancing their reasoning capabilities (jaech2024openai; guo2025deepseek; luo2025deepscaler; hu2025open; zeng2025simplerl). However, the effectiveness of RFT is highly sensitive to task selection (parashar2025curriculum; shen2025skywork; zhu2025eduflow; wen2025structured; li2025temple). Naively training on a static, uniformly sampled dataset is inefficient: the model spends excessive computation on tasks that are either already mastered (too easy) or beyond reach (too hard) (yu2025dapo; bae2025online; chen2025self). This inefficiency not only inflates training costs but also destabilizes optimization by reducing the effective batch size. The central challenge, therefore, is to dynamically select tasks of “just right” difficulty to maximize learning efficiency as the model’s capability evolves.
Existing methods to this challenge face several limitations. Offline task selection (parashar2025curriculum; shen2025skywork; zhu2025eduflow; wen2025structured; li2025temple), which pre-schedules tasks from easy to hard, is too rigid and does not adapt to the evolving trajectory of the model. In response, a few online selection methods have been proposed, aiming to adaptively choose tasks based on model’s current capability. Core challenge of these methods lies in the tradeoff between the computational cost of collecting information and the accuracy of the resulting performance estimates. We argue that existing solutions are not sufficiently efficient: some expend excessive computation on information gathering, undermining efficiency, while others fail to fully exploit collected information, leading to suboptimal selection. On one hand, oversampling-based methods (yu2025dapo; bae2025online) find suitable tasks by rolling out oversized batches, introducing substantial extra cost. On the other hand, non-oversampling approaches typically rely on a single source of information—either leveraging historical evaluations as explicit evidence (chen2025self; qu2025can) or exploiting inter-task correlations as implicit evidence (sun2025improving). Our empirical results reveal a clear complementarity: explicit evidence provides stable and accurate task-difficulty estimates but suffers from a slow warm-up when historical evaluations are scarce in early training, whereas implicit evidence quickly guides early-stage selection yet becomes less reliable in later stages. These findings indicate that relying solely on one type of evidence leaves information underutilized and leads to suboptimal task selection. Therefore, a principled framework to fuse these complementary evidence sources is essential for robust and efficient online task selection.
In this work, we introduce BOTS, the first unified and extensible framework for Bayesian Online Task Selection in LLM reinforcement finetuning. BOTS recasts online task selection as a principled Bayesian inference problem over the model’s evolving capabilities. By doing so, it naturally addresses the core challenges of non-stationarity and partial observability, featuring with three key design elements: (1) Bayesian foundation: Grounded in Bayesian inference, the framework naturally adapts to the evolving capability of the model, allowing task difficulty to be continuously re-estimated. (2) Integration of two evidence sources: Tunable update rules jointly incorporate explicit evidence from direct evaluations and implicit evidence inferred from related tasks, leveraging their complementary strengths. (3)Thompson sampling: Task selection is guided by posterior sampling, ensuring a principled balance between exploration and exploitation.
For implicit evidence, we further instantiate the framework with an extremely efficient interpolation-based plugin that estimates the difficulty of unevaluated tasks without additional rollouts, making the overhead negligible. We demonstrate empirically, across diverse domains and model scales, that our method significantly improves data efficiency and model performance over baselines and ablations, offering a practical, effective, and extensible solution for online task selection in RFT.
Figure 1: Overview of the BOTS framework. BOTS operates in a continuous loop of task selection, model training, and posterior updating. (1) Selection: Thompson sampling from the posterior beliefs selects a batch of tasks whose estimated success probabilities are near a target difficulty (e.g., $p^{*}=0.5$). (2) Training & Evidence Collection: The LLM is finetuned, yielding direct success/failure counts (explicit evidence) for the selected batch. For unselected tasks, predicted counts (implicit evidence) are produced by a plug-in; in Section 3.3, we introduce an ultra-lightweight interpolation-based variant with negligible overhead. (3) Posterior Updating: Explicit and implicit evidence are fused using our generalized Bayesian update rule (Section 3.2).
2 Related Works
The impact of task difficulty on RFT of LLMs has become an active research topic. Inspired by the seminal idea of curriculum learning (bengio2009curriculum), researchers have proposed various strategies for selecting appropriate tasks in LLM RFT. For instance, parashar2025curriculum; shen2025skywork advocate scheduling tasks from easy to hard, enabling LLMs to gradually acquire reasoning skills. Similar ideas have been extended to multi-modal LLMs (zhu2025eduflow; wen2025structured; li2025temple), though these methods mainly focus on offline task selection following an easy-to-hard trajectory.
More recently, online selection strategies have emerged, often targeting tasks of moderate difficulty. One line of work adopts sampling-based task filtering, where tasks with consistently trivial rewards (all zeros or ones) are considered uninformative and are down-weighted or filtered (yu2025dapo; bae2025online). While effective, these methods require additional rollouts, incurring non-trivial overhead. To avoid extra rollouts, several works attempt to predict task passing rates without direct rollouts. chen2025self formulate task selection as a non-stationary multi-armed bandit problem, treating each problem category (e.g., difficulty level or type) as an arm and using absolute advantage as a reward proxy, with posterior estimation based on historical outcomes. qu2025can extend this framework to the task level. Although these methods eliminate extra rollouts, they rely solely on direct evaluations and overlook cross-task relationships. In contrast, sun2025improving propose evaluating a small set of reference tasks and predicting the passing rates of others using an attention-inspired kernel over embeddings. However, this approach still requires additional rollouts for the reference set and discards historical evaluation information.
More comprehensive discussion and comparison on related works are provided in Appendix C.
3 Bayesian Online Task Selection
3.1 Preliminaries: Modeling Task Difficulty
A task $\mathcal{T}=(Q,R)$ is defined as a tuple consisting of a query $Q$, expressed in natural language, and a reward function $R$ that maps any natural language response $O$ to a binary reward $R(Q,O)\in\{0,1\}$, which is common in domains like math, coding such that $1$ indicates correct and 0 indicates incorrect. Consider RFT of a parameterized language model $\mathcal{M}_{\theta}$, which maps a query $Q$ to a response $O$, on a set of $N$ tasks $\{\mathcal{T}^{k}\}_{k=1}^{N}$. The binary reward obtained by executing the model on a task $\mathcal{T}$ follows a Bernoulli distribution $\texttt{Bernoulli}(p_{\theta,\mathcal{T}})$, where $p_{\theta,\mathcal{T}}=\mathbb{E}_{o\sim\mathcal{M}(\cdot\mid\mathcal{T};\theta)}\,R(O,\mathcal{T})$ denotes the model’s success probability on $\mathcal{T}$. With a slight abuse of notation, we denote the reward distribution for a given model and task as $R(\cdot|\mathcal{T};\theta):=\texttt{Bernoulli}(p_{\theta,\mathcal{T}})$. Since we focus on online task selection over a fixed set of tasks, we simplify the notation by letting $p_{t}^{k}$ denote $p_{\theta_{t},\mathcal{T}^{k}}$ and $R_{t}^{k}$ denote $R(\cdot|\mathcal{T}^{k};\theta_{t})$, whenever the context is clear. All notations are summarized in Table 5.
3.2 Core Mechanism: Fusing Evidence in a Unified Posterior
Our goal is to estimate the success probability $p_{t}^{k}$ of the online-adapted model on $k$-th task $\mathcal{T}^{k}$. For efficiency, direct evaluations are only performed after a task is selected. As statistical evidence, at time step $t$, we obtain online samples $r^{k}_{1:n}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}R_{t}^{k}(\cdot)$ for each selected task $\mathcal{T}^{k}$ in the training batch $\mathcal{B}_{t}$, where $n$ corresponds to the number of rollouts per task.
A natural way to model the estimation is via a Beta distribution, $\texttt{Beta}(\alpha_{t}^{k},\beta_{t}^{k})$, where the posterior parameters $\alpha_{t}^{k}$ and $\beta_{t}^{k}$ represent the accumulated counts of successes and failures, respectively, for model $\theta_{t}$ on task $\mathcal{T}^{k}$. The problem then reduces to designing online adaptation rules for $\alpha_{t}^{k}$ and $\beta_{t}^{k}$. We propose the following online adaptation rules: Given a batch of direct evaluation results $\mathcal{B}_{t}=\{(\mathcal{T}_{\mathcal{B}_{t}[i]},r_{1:n}^{\mathcal{B}_{t}[i]})\}_{i=1}^{|\mathcal{B}_{t}|}$, we define the adaptation rules as
$\alpha_{t+1}^{k}=(1-\lambda)\alpha_{t}^{k}+\lambda\alpha_{0}^{k}+(1-\rho)\,s_{t}^{k}+\rho\,\tilde{s}_{t}^{k},\qquad\beta_{t+1}^{k}=(1-\lambda)\beta_{t}^{k}+\lambda\beta_{0}^{k}+(1-\rho)\,f_{t}^{k}+\rho\,\tilde{f}_{t}^{k},$ (1)
where $\alpha_{0},\beta_{0}$ denote the prior parameter set for the Beta distribution, and the coefficient $\lambda\in[0,1]$ discounts historical information;
$s_{t}^{k}=\sum_{k^{\prime}\in\mathcal{B}_{t}}\mathbb{I}[k^{\prime}=k]\sum_{i=1}^{n}r_{i}^{k^{\prime}},\qquad f_{t}^{k}=\sum_{k^{\prime}\in\mathcal{B}_{t}}\mathbb{I}[k^{\prime}=k]\sum_{i=1}^{n}(1-r_{i}^{k^{\prime}})$ (2)
denote the explicit success and failure counts from direct evaluations, by slightly abusing the notation $k\in\mathcal{B}_{t}$ to represent task $\mathcal{T}^{k}$ received direct evaluation at time step $t$. Notice when direct evaluation results are not available ($k\notin\mathcal{B}_{t}$), $s_{t}^{k}=f_{t}^{k}=0$; and
$\tilde{s}_{t}^{k}=s_{t}^{k}+\mathbb{I}[k\notin\mathcal{B}_{t}]\,\tilde{p}(k,\mathcal{B}_{t})\,n,\qquad\tilde{f}_{t}^{k}=f_{t}^{k}+\mathbb{I}[k\notin\mathcal{B}_{t}]\,(1-\tilde{p}(k,\mathcal{B}_{t}))\,n.$ (3)
Here, $\rho\in[0,1]$ balances the contributions of explicit and implicit evidence, $\tilde{s}_{t}^{k}$ and $\tilde{f}_{t}^{k}$ coincide with $s_{t}^{k}$ and $f_{t}^{k}$ when direct evaluation results are available for task $\mathcal{T}_{k}$, and otherwise represent the pseudo success and failure counts. These are derived from an estimator $\tilde{p}(k,\mathcal{B}_{t})$, which uses inter-task relationships to infer difficulty for tasks not present in the current evaluation batch $\mathcal{B}_{t}$. Our framework places no restrictions on the specific form of $\tilde{p}(k,\mathcal{B}_{t})$, while in Sec. 3.3, we introduce a lightweight interpolation-based instance to produce the pseudo counts. Additionally, to manage the equivalent total sample size—and hence the uncertainty of the estimate—the pseudo sample size is ensured to satisfy $\tilde{s}_{t}^{k}+\tilde{f}_{t}^{k}=n$.
The following proposition indicates that the update in Equation (1) preserves the Beta family as the (generalized) posterior for a Bernoulli parameter under a tempered/prior-mixing update.
{restatable}
propositionpropositionone Let $p\in(0,1)$ be the Bernoulli success probability at time $t$. Suppose the current belief is $\pi_{t}(p)=\mathrm{Beta}(p\mid\alpha_{t},\beta_{t})$, and let $\pi_{0}(p)=\mathrm{Beta}(p\mid\alpha_{0},\beta_{0})$ be a base prior. Given counts $(s_{t},f_{t})$ and pseudo counts $(\tilde{s}_{t},\tilde{f}_{t})$ with $s_{t},f_{t},\tilde{s}_{t},\tilde{f}_{t}\geq 0$, define the generalized-Bayes update
[…]
with $\lambda\in(0,1)$ and $\rho\in[0,1]$. Then $\pi_{t+1}$ is exactly $\mathrm{Beta}(\alpha_{t+1},\beta_{t+1})$ with
$\alpha_{t+1}=(1-\lambda)\alpha_{t}+\lambda\alpha_{0}+(1-\rho)s_{t}+\rho\tilde{s}_{t},\qquad\beta_{t+1}=(1-\lambda)\beta_{t}+\lambda\beta_{0}+(1-\rho)f_{t}+\rho\tilde{f}_{t}.$
The proof is placed in Appendix D.1.
3.3 Ultra-Light Interpolation Plug-in for Implicit Evidence
Given a batch of online evaluation results $\mathcal{B}_{t}=\{(\mathcal{T}_{\mathcal{B}_{t}[i]},r_{1:n}^{\mathcal{B}_{t}[i]})\}_{i=1}^{|\mathcal{B}_{t}|}$, we aim to estimate the passing rate $p_{t}^{k}$ for any task $\mathcal{T}_{k}$ using an estimator $\tilde{p}(\mathcal{B}_{t},k)$. In this work, we adopt an ultra-lightweight interpolation-based estimator to minimize additional computational overhead for online task selection. Notably, the adaptation rules in Sec. 3.2 place no restrictions on the specific form of $\tilde{p}(\mathcal{B}_{t},k)$.
Assume that for each task $\mathcal{T}^{k}$, we have empirical success rates $\bar{p}_{w}^{k}$ and $\bar{p}_{s}^{k}$ from two reference models of distinct capability (weak vs. strong). Define the average empirical success rates of the current, weak, and strong models on $\mathcal{B}_{t}$ as
$\bar{p}_{t}^{\mathrm{ref}}(\mathcal{B}_{t}):=\frac{1}{|\mathcal{B}_{t}|}\sum_{k\in\mathcal{B}_{t}}\frac{1}{n}\sum_{j=1}^{n}r_{j}^{k},\quad\bar{p}_{w}^{\mathrm{ref}}(\mathcal{B}_{t}):=\frac{1}{|\mathcal{B}_{t}|}\sum_{k\in\mathcal{B}_{t}}\bar{p}_{w}^{k},\quad\bar{p}_{s}^{\mathrm{ref}}(\mathcal{B}_{t}):=\frac{1}{|\mathcal{B}_{t}|}\sum_{k\in\mathcal{B}_{t}}\bar{p}_{s}^{k}.$
We estimate the relative capability coefficient of the current model as $\mu_{t}(\mathcal{B}_{t})=\big(\bar{p}_{t}^{\mathrm{ref}}(\mathcal{B}_{t})-\bar{p}_{w}^{\mathrm{ref}}(\mathcal{B}_{t})\big)/\ \big(\bar{p}_{s}^{\mathrm{ref}}(\mathcal{B}_{t})-\bar{p}_{w}^{\mathrm{ref}}(\mathcal{B}_{t})\big),$ which locates the current model between the weak and strong reference models on the batch $\mathcal{B}_{t}$ (we assume $\bar{p}_{s}^{\mathrm{ref}}>\bar{p}_{w}^{\mathrm{ref}}$; otherwise one may add a small $\varepsilon$ to the denominator). To reduce variance from stochastic rollouts, we maintain a momentum version of the coefficient $\tilde{\mu}_{t}=\gamma\tilde{\mu}_{t-1}+(1-\gamma)\mu_{t}$.
Finally, the passing rate of the current model on task $\mathcal{T}^{k}$ is obtained via linear interpolation between the weak and strong references, followed by clipping to $[0,1]$:
$\tilde{p}(\mathcal{B}_{t},k)=\mathrm{clip}\,\Big(\tilde{\mu}_{t}(\mathcal{B}_{t})\,\bar{p}_{s}^{k}+\bigl(1-\tilde{\mu}_{t}(\mathcal{B}_{t})\bigr)\,\bar{p}_{w}^{k},\;0,\;1\Big).$ (5)
Computational Cost. The primary cost of our approach arises from evaluating the reference models for all tasks, which requires extensive rollouts. Fortunately, benchmarking standard base models has already become routine in RL dataset construction—for example, cheng2025revisiting report Qwen2.5-7B-Instruct (Yang2024Qwen25TR) and Qwen3-30B-A3B (yang2025qwen3) performance as meta-tags for filtering. Given these pre-computed references, the additional online overhead of our interpolation-based estimator is negligible compared to training-time rollouts (see Appendix F.1 for empirical results). This efficiency follows from two properties: (i) no extra rollouts are needed during training, and (ii) both updates and predictions reduce to lightweight vector operations.
3.4 Thompson Sampling for Task Selection
Having established a task difficulty posterior $\mathrm{Beta}(p_{t}^{k}\mid\alpha_{t}^{k},\beta_{t}^{k})$ over the success probability of each task $\mathcal{T}^{k}$, we now turn to the crucial step of selecting tasks for the next training batch.
The first question is: at what difficulty level does the current model benefit most from training? Prior works (chen2025self; sun2025improving) show that, under binary rewards, tasks with success probability around $0.5$ are most informative for learning as they lead to gradients with larger expected magnitude than tasks with success probability close to 0 or $1$. We define the utility of a task as the absolute deviation of its posterior mean from a target success probability $p^{*}\in(0,1)$, with $p^{*}=0.5$ as the canonical choice.
Given this target success probability, the problem of online task selection naturally reduces to a non-stationary bandit problem. The central challenge is the tradeoff between exploitation and exploration: the model must decide whether to select tasks with high-confidence estimates close to the target rate in order to maximize immediate utility, or to select tasks with high uncertainty to gather information that may improve future decisions. A purely exploitative strategy might select tasks whose posterior mean $\tilde{p}_{t}^{k}$ is closest to $p^{*}$, but this risks overlooking tasks whose difficulty is currently uncertain but potentially optimal. To naturally balance exploration and exploitation, we employ Thompson Sampling (thompson1933likelihood), a strategy renowned for both its empirical effectiveness and theoretical guarantees. More specifically, at each selection step $t$, we perform the following: (1) Posterior Sampling: Draw a sample of the passing rate from its current posterior distribution $\hat{p}_{k}\sim\mathrm{Beta}(\alpha_{t}^{k},\beta_{t}^{k})$ for each task $\mathcal{T}^{k}$ in the pool. (2) Selection: Select tasks with the highest estimated utilities $\{\hat{u}_{k}:=|\hat{p}_{k}-p^{*}|\}$ to form the training batch $\mathcal{B}_{t+1}$. This procedure elegantly prioritizes tasks that are likely to be near the target difficulty, while the inherent variance in sampling from the posterior ensures that tasks with higher uncertainty are naturally explored.
Impact of $\lambda,\rho$ on Task Selection. Proposition 3.4 characterizes how the hyperparameters $\lambda$ and $\rho$ jointly determine the effective sample size $n_{t}=\alpha_{t}+\beta_{t}$ of each task’s Beta posterior, thereby controlling the confidence of the estimated probability of success.
{restatable}
propositionpropositiontwo Let $n_{t}:=\alpha_{t}+\beta_{t}$. Suppose the updates follow Equation (1)–(3) with $\lambda\in(0,1)$, $\rho\in[0,1]$, we have
$\liminf_{t\to\infty}n_{t}\;=\;n_{0}+\tfrac{\rho}{\lambda}n,\qquad\limsup_{t\to\infty}n_{t}\;=\;n_{0}+\tfrac{1}{\lambda}n.$
The proof is given in Appendix D.2. Specifically, $\lambda$ controls the overall scale of the effective sample size: smaller values accelerate forgetting and increase posterior uncertainty, promoting exploration, while larger values stabilize the posterior and favor exploitation. Meanwhile, $\rho$ sets the lower bound of effective counts for unevaluated tasks: smaller values maintain high uncertainty to encourage exploration, whereas larger values reduce uncertainty and bias selection toward exploitation. Together, $\lambda$ and $\rho$ provide complementary knobs for balancing exploration and exploitation in online task selection. A more detailed discussion is placed in Appendix E.1.
4 Experiments
We begin with Section 4.1, which introduces datasets, reinforcement finetuning protocols, evaluation metrics, and computational cost. Section 4.2 and Section 4.3 analyze the effects of mixing evidence ($\rho$) and forgetting/adaptivity ($\lambda$) on task selection respectively. Section 4.4 then compares BOTS with competitive baselines across model scales and domains, while Section 4.5 summarizes additional experiments provided in the Appendix F.
4.1 Setups
Dataset.  We conduct experiments on GURU (cheng2025revisiting), a well-curated cross-domain RL dataset. Each subset is deduplicated, verified, and filtered. We use its math, code, and logic subsets (excluding the Zebra Puzzle due to its non-binary reward). Detailed information about the used datasets is provided in Appendix B.1 (for training) and Appendix B.2 (for evaluation).
RFT Setting.  We adopt GRPO (grpo2024), Qwen2.5-1.5B-Instruct and Qwen2.5-7B. Key hyperparameters include a learning rate of 1e-6, 16 rollouts per task, and a temperature of $1.0$. Comprehensive training details and used RL algorithm are provided in Appendix B.3 and B.5.
Evaluation.  We report the following metrics to evaluate our framework and its ablations, with formal definitions given in Appendix B.4.
• Effective Task Ratio (ETR).  It evaluates tasks selection. The fraction of sampled tasks whose empirical success rate, estimated from $n=16$ independent rollouts, falls strictly within the $(0,1)$ range. A higher ETR indicates a more efficient task selection that successfully filters out tasks that are either already mastered ($p=1$) or currently unsolvable ($p=0$).
• Time-to-Baseline (TTB).  It measures training acceleration relative to the random baseline to achieve a specific performance. Let the baseline start from performance $P_{\text{init}}$ and reach the best performance $P_{\text{best}}$ within the training window. For a target fraction $\tau\in\{50\%,75\%,100\%\}$, we define the target performance as $P_{\tau}=P_{\text{init}}+\tau\cdot(P_{\text{best}}-P_{\text{init}}).$ TTB($\tau$) is the ratio of steps required by a method to reach $P_{\tau}$ compared to the baseline. For example, if the baseline starts at $0.1$ and reaches $0.3$ by step 100, then the 50% target is $0.2$. If the baseline first reaches $0.2$ at step 40 while another method reaches it at step 30, then $\text{TTB}(50\%)=30/40=0.75$. By definition, the baseline has $\text{TTB}=1$; smaller values indicate greater acceleration.
• Best-so-far (BSF).  It measures the performance gain relative to the random baseline under a fixed budget. Within specific ratio (25%, 50%, 100%) of total training steps, BSF is the ratio between a method’s best-so-far performance and the baseline’s best-so-far performance. For example, at step 50 (total steps 100), if the baseline’s best is 0.4 and a method’s best is 0.6, then BSF(50%) = $0.6/0.4=1.5$. The baseline always has BSF = 1; larger values indicate greater gains.
Computational Overhead. We note that BOTS introduces negligible additional computation, with overhead measured at $\leq 0.2\%$ of total training time (see analysis in Section 3.3 and empirical results in Appendix F.1). Thus, in the main results, we report TTB and BSF, in terms of training steps rather than wall-clock time, as their difference is practically insignificant.
4.2 Fusing Two sources of Evidence: Analyzing the Impact of $\rho$
Recall as defined in Equation 1, $\rho$ controls the relative weights of explicit and implicit evidence. A smaller $\rho$ places greater emphasis on explicit evidence, whereas a larger $\rho$ increases the contribution of implicit evidence. Moreover, according to Proposition 3.4 and the analysis in Section 3.4, $\rho$ determines the limit inferior of the equivalent counts in the Beta posterior. Intuitively, reducing $\rho$ increases the overall randomness of posterior sampling, while enlarging $\rho$ stabilizes the estimates by anchoring them more strongly to implicit evidence.
We investigate the role of $\rho$ in task selection by varying $\rho\in\{0.0,0.05,0.1,0.2,0.5,1.0\}$, while keeping other hyperparameters fixed at their default values ($\lambda=0.1$, posterior sampling enabled).
Figure 2: Qwen2.5-1.5B-Instruct on Math. Ratio of sampled training tasks (measured over 16 rollouts) with passing rates: strictly between 0 and 1, strictly greater than 0, and strictly less than 1, along with aggregated performance (MATH500 and AIME24), plotted against training steps.
Benchmark MATH500 AIME24 Math Aggregated Performance
Metric TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$)
Target Fraction 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100%
Random 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
$\lambda=0.1,\rho=0.0$ 1.10 0.73 0.88 1.00 0.96 1.03 0.57 0.92 - 1.00 1.75 0.92 1.13 0.73 0.99 1.02 0.98 1.01
$\mathbf{\lambda=0.1,\rho=0.05}$ 1.02 0.56 0.60 1.08 1.06 1.05 0.50 0.51 0.68 1.00 2.50 1.17 1.07 0.58 0.60 1.10 1.05 1.03
$\mathbf{\lambda=0.1,\rho=0.1}$ 0.89 0.49 0.57 1.13 1.05 1.05 0.51 1.00 1.00 1.25 1.75 1.00 0.89 0.56 0.64 1.12 1.07 1.05
$\mathbf{\lambda=0.1,\rho=0.2}$ 0.99 0.68 0.68 1.10 1.10 1.06 0.96 1.06 - 1.25 1.50 0.92 0.99 0.74 0.80 1.13 1.08 1.08
$\lambda=0.1,\rho=0.5$ 0.93 0.64 0.67 1.10 1.10 1.03 0.63 1.11 1.12 0.75 2.25 1.00 0.93 0.66 1.00 1.10 1.11 1.01
$\lambda=0.1,\rho=1.0$ 0.92 0.61 0.95 1.04 1.08 1.02 0.42 0.71 - 1.25 2.25 0.92 0.90 0.61 0.86 1.05 1.06 1.01
Table 1: TTB and BSF evaluated on MATH500, AIME24, and aggregated performance with Qwen2.5-1.5B-Instruct. For TTB, notation “-” indicates that the target performance is never achieved within the evaluation window. The best and second best results are marked accordingly.
Analysis and Takeaways. Our central finding is that implicit evidence is crucial for rapid cold-starts, while explicit evidence is vital for long-term accuracy. A principled fusion of both is key to effective task selection. Relying solely on implicit evidence ($\rho=1$) leads to error accumulation, while ignoring it ($\rho=0$) suffers from severe data sparsity in early training, especially on large-scale datasets. This conclusion is supported by the following observations from Figure 2 and Table 1:
(1) Implicit evidence provides an essential early boost. In the initial training phase, all settings with $\rho>0$ demonstrate a sharp increase in the Effective Task Ratio (ETR) over the random baseline, primarily by filtering out unsolvable tasks ($p=0$). In contrast, the $\rho=0$ setting, which relies solely on sparse explicit feedback, behaves almost identically to random sampling. This highlights the critical role of implicit evidence in overcoming the cold-start problem.
(2) Over-reliance on implicit evidence degrades long-term performance. As training progresses, settings with a large $\rho$ (e.g., 0.5, 1.0) show a declining ETR, as the implicit evidence cannot perfectly track the learning model’s fine-grained progress. It often fails to recognize when tasks are fully mastered ($p=1$), leading to their wasteful selection. This inaccuracy ultimately harms late-stage performance, as seen in the aggregated TTB and BSF metrics where large $\rho$ values underperform.
(3) A small positive $\rho$ achieves the best of both worlds. A $\rho\in\{0.05,0.1,0.2\}$ strikes an optimal balance. It leverages implicit evidence for initial acceleration while allowing more accurate, accumulating explicit evidence to dominate in the long run. This fusion leads to robust performance gains throughout entire training process, achieving best overall TTB and BSF scores.
4.3 Forgetting and Adaptivity: Analyzing the Impact of $\lambda$
Recall as defined in Equation 1, $\lambda$ controls the discounting of historical information in the Beta posterior. A smaller $\lambda$ places more emphasis on recent evaluations, improving adaptability but increasing the variance of the estimates. Conversely, a larger $\lambda$ gives greater weight to accumulated history, yielding more stable but less responsive estimates. Furthermore, Proposition 3.4 shows that $\lambda$ determines the scale of the equivalent counts in the Beta posterior. Intuitively, reducing $\lambda$ accelerates the decay of equivalent counts, increasing the randomness of posterior sampling, whereas enlarging $\lambda$ slows the decay and stabilizes the estimates.
We investigate the role of $\lambda$ in task selection by varying $\lambda\in\{0.0,0.05,0.1,0.2,0.5,1.0\}$, while keeping other hyperparameters fixed at their default values ($\rho=0.1$, posterior sampling enabled).
Figure 3: Qwen2.5-1.5B-Instruct on Math. Ratio of sampled training tasks (measured over 16 rollouts) with passing rates: strictly between 0 and 1 (L1), strictly greater than 0 (L2), and strictly less than 1 (L3), along with MATH500 Accuracy (avg@1), plotted against training steps.
Benchmark MATH500 AIME24 Math Aggregated Performance
Metric TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$)
Target Fraction 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100%
Random 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
$\lambda=0.0,\rho=0.1$ 0.88 0.57 0.62 1.09 1.11 1.03 0.77 - - 1.25 1.50 0.75 0.87 0.58 0.71 1.11 1.08 1.04
$\mathbf{\lambda=0.05,\rho=0.1}$ 0.92 0.45 0.75 1.12 1.07 1.02 0.64 1.11 1.16 1.25 1.75 1.08 0.94 0.50 0.76 1.15 1.07 1.05
$\mathbf{\lambda=0.1,\rho=0.1}$ 0.89 0.49 0.57 1.13 1.05 1.05 0.51 1.00 1.00 1.25 1.75 1.00 0.89 0.56 0.64 1.12 1.07 1.05
$\mathbf{\lambda=0.2,\rho=0.1}$ 0.92 0.56 0.59 1.08 1.08 1.06 0.63 - - 0.75 2.00 0.75 0.95 0.58 0.66 1.07 1.11 1.06
$\lambda=0.5,\rho=0.1$ 1.11 0.53 0.77 1.11 1.07 1.03 0.51 0.89 - 1.25 1.75 0.83 1.14 0.74 0.86 1.12 1.06 1.03
$\lambda=1.0,\rho=0.1$ 1.15 0.68 0.80 1.04 1.04 1.03 0.60 0.63 - 1.00 2.50 0.83 1.14 0.68 1.00 1.04 1.09 1.02
Table 2: TTB and BSF evaluated on MATH500, AIME24, and aggregated performance with Qwen2.5-1.5B-Instruct. For TTB, notation “-” indicates that the target performance is never achieved within the evaluation window. The best and second best results are marked accordingly.
Analysis and Takeaways. The hyperparameter $\lambda$ governs the framework’s adaptivity to the non-stationary learning process. Our key takeaway is that task selection must be adaptive, and moderate forgetting is essential for long-term performance. An overly long memory ($\lambda\to 0$) prevents the system from recognizing when a task has been mastered, while an overly short memory ($\lambda\to 1$) leads to unstable estimates. This conclusion is supported by the following observations from Figure 3 and Table 2:
(1) Small $\lambda$ struggles with mastered tasks: Settings with small $\lambda$ (e.g., 0.0, 0.05) exhibit a declining ETR in the mid-to-late training phase. This is because their long memory makes them slow to update the difficulty of tasks that the model has recently mastered (from $p<1$ to $p=1$), causing them to be wastefully selected. This ultimately limits late-stage performance gains.
(2) Large $\lambda$ lacks stability: Conversely, large $\lambda$ values (e.g., 0.5, 1.0) correspond to rapid forgetting. This makes the posterior estimates highly volatile and overly dependent on the most recent (and potentially noisy) observations. As a result, they fail to build stable beliefs about task difficulties, leading to only marginal improvements over the random baseline in the early training stages.
(3) Moderate $\lambda$ balances stability and adaptivity: A moderate $\lambda\in\{0.05,0.1,0.2\}$ strikes a balance. It provides enough memory to form stable difficulty estimates while being adaptive enough to track the model’s evolving capabilities. This leads to strong and consistent performance gains across the entire training process, achieving the best TTB and BSF results.
4.4 Performance Comparison Across Models and Domains
Domain Math Code Logic
Metric TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$)
Target Fraction 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100%
Random 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Offline 0.96 0.51 0.78 1.16 1.12 1.01 0.68 1.03 - 1.09 1.00 0.99 1.23 1.36 - 0.67 0.78 0.89
BOTS-MoPPS 1.78 1.32 - 0.90 0.90 0.99 0.78 0.96 1.09 1.09 1.02 1.02 0.87 1.04 - 1.13 0.96 0.96
BOTS-DOTS 0.83 0.75 0.81 1.11 1.09 1.08 0.67 0.65 0.79 1.17 1.09 1.02 0.66 1.32 - 1.28 0.92 0.93
BOTS 0.89 0.56 0.64 1.12 1.07 1.05 0.58 0.90 0.77 1.17 1.08 1.03 0.85 0.94 1.05 1.19 1.01 1.00
[…]
Domain Math Code Logic
Metric TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$)
Target Fraction 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100%
Random 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Offline 0.91 0.73 0.76 0.99 1.00 1.05 0.97 1.24 - 0.99 0.99 0.99 1.06 1.23 0.95 0.96 0.95 1.04
BOTS-MoPPS 1.07 1.15 0.70 0.94 1.04 1.04 0.84 0.69 0.84 1.04 1.02 1.00 0.75 0.92 0.69 1.07 1.00 1.05
BOTS-DOTS 0.83 0.80 0.61 0.98 1.02 1.03 0.86 0.96 0.76 1.02 1.04 1.02 0.79 0.91 0.91 1.02 0.95 1.03
BOTS 0.86 0.77 0.63 0.99 1.04 1.04 0.82 0.79 1.06 1.04 1.01 1.00 0.79 0.78 0.50 1.03 1.01 1.00
[…]
We conduct extended experiments across model sizes (1.5B and 7B) and task domains (math, code, logic). We compare BOTS under the default setting ($\lambda=0.1,\rho=0.1$, posterior sampling enabled) against four baselines. Two are out-of-framework baselines: (i) Random, where tasks are uniformly sampled, and (ii) Offline, where tasks are ranked from easy to hard based on success probabilities of Qwen2.5-7B-Instruct and Qwen3-30B-A3B (for tie-breaking). Two are within-framework baselines: (iii) BOTS-MoPPS, with $\lambda=0.0,\rho=0.0$ and posterior sampling enabled, which reduces our framework to MoPPS (qu2025can) and thus enables direct evaluation of a purely explicit-evidence strategy; and (iv) BOTS-DOTS, with $\lambda=1.0,\rho=1.0$ and posterior sampling disabled, serving as a proxy inspired by DOTS (sun2025improving), which evaluates the long-term effectiveness of relying almost entirely on implicit evidence without corrective feedback. Implementation details and further discussion of these baselines are provided in Appendix B.6.
Analysis and Takeaways. The default BOTS setting consistently outperforms both out-of-framework and within-framework baselines, validating the principle of fusing explicit and implicit evidence. Moreover, BOTS-DOTS emerges as the strongest baseline, confirming that our interpolation-based implicit evidence provides useful guidance for task selection. These conclusions are supported by the following observations from Table 3 and Table 4:
(1) BOTS achieves the best overall performance. For the 1.5B model, BOTS obtains 8 first-place and 9 second-place finishes out of 18 reported metrics, with a notable 36% acceleration (TTB(100%) = 0.64) in the math domain. For the 7B model, BOTS secures 6 first-place and 8 second-place finishes, including a remarkable 50% acceleration (TTB(100%) = 0.50) in the logic domain.
(2) BOTS-DOTS ranks second. For the 1.5B model, BOTS-DOTS achieves 6 first-place and 4 second-place finishes. For the 7B model, it achieves 5 first-place and 1 second-place finishes, outperforming the remaining baselines.
In summary, these results demonstrate that BOTS not only delivers consistent gains across different domains and model scales, but also achieves substantial acceleration in training efficiency. The superiority of BOTS over both BOTS-MoPPS and BOTS-DOTS highlights the necessity of combining explicit and implicit evidence, while the strength of BOTS-DOTS underscores the practical value of our interpolation-based implicit evidence. Together, these findings establish BOTS as a principled, effective, and scalable solution for online task selection in LLM RFT.
4.5 Overview of Additional Experiments
We provide extended experiments in Appendix F for a deeper understanding of BOTS: (1) A wall-clock breakdown (Appendix F.1) shows task selection adds less than 0.2% overhead. (2) Offline evaluation of the interpolator (Appendix F.2) confirms it effectively tracks task difficulty. (3) An ablation on Thompson sampling (Appendix F.3) shows posterior sampling yields more stable selection. (4) A fine-grained analysis of selected task dynamics (Appendix F.4) illustrates how BOTS shifts computation away from trivial ($p=1$) and impossible ($p=0$) tasks toward the informative mid-difficulty region.
5 Conclusion and Discussion
We introduced BOTS, a unified Bayesian framework for online task selection in LLM reinforcement finetuning. BOTS formulates task difficulty estimation as Bayesian belief updating, fusing two complementary evidence sources: stable but sparse explicit evidence from direct rollouts and dense but less precise implicit evidence from inter-task relationships. Instantiated with a lightweight interpolator and Thompson sampling, BOTS achieves adaptive and efficient online data selection with negligible overhead. Experiments show consistent gains in data efficiency and model performance across domains and model scales, surpassing methods that rely on a single evidence source. We envision BOTS as a practical foundation for dynamic, model-aware data selection, advancing efficient and effective LLM training.
This work opens several promising directions. First, we focused mainly on binary-reward tasks; extending and validating BOTS in non-binary reward settings is an important next step. Second, BOTS currently uses fixed update rules, though our results show that different values of $\lambda$ and $\rho$ work best at different training stages. Developing adaptive update strategies that adjust to training dynamics would further improve robustness. Third, while our lightweight interpolator efficiently provides implicit evidence, designing stronger plug-in alternatives and systematically studying the trade-off between predictive accuracy and computational cost remain open challenges. A more detailed discussion of these directions is provided in Appendix E.4.
Appendix
Appendix A Notation Summary
For ease of reading and reference, we present the symbols used in this paper in Table 5.
Symbol Description
General & Problem Setup
$\mathcal{T}_{k},\mathcal{T}$ The $k$-th task, or a generic task.
$Q,O$ A query (input) and an outcome (output) of a task.
$R(\cdot)$ A binary reward function, mapping an outcome to $\{0,1\}$.
$\mathcal{M}_{\theta}$ The language model parameterized by $\theta$.
$p_{\theta,\mathcal{T}}$ Success probability of model $\mathcal{M}_{\theta}$ on task $\mathcal{T}$.
$t$ Index for the training step or iteration.
$p_{t}^{k}$ Shorthand for $p_{\theta_{t},\mathcal{T}_{k}}$, success probability on task $k$ at step $t$.
$R_{t}^{k}$ Shorthand for the reward distribution (Bernoulli) of task $k$ at step $t$.
NN Total number of tasks in the pool.
$n$ Number of rollouts per task in a batch.
$\mathcal{B}_{t}$ The batch of tasks selected for training at step $t$.
Bayesian Estimation Framework
$\alpha_{t}^{k},\beta_{t}^{k}$ Parameters of the Beta posterior for task $k$ at step $t$.
$\alpha_{0}^{k},\beta_{0}^{k}$ Parameters of the base prior Beta distribution for task $k$.
$n_{t}^{k}$ Equivalent sample size for task $k$ at step $t$, $n_{t}^{k}=\alpha_{t}^{k}+\beta_{t}^{k}$.
$n_{0}^{k}$ Equivalent sample size of the prior for task $k$, $n_{0}^{k}=\alpha_{0}^{k}+\beta_{0}^{k}$.
$\lambda$ Forgetting factor in $[0,1]$ for historical information.
$\rho$ Balance coefficient in $[0,1]$ for explicit vs. implicit evidence.
$r_{i}^{k}$ The $i$-th binary reward obtained for task $k$.
$s_{t}^{k},f_{t}^{k}$ Explicit success and failure counts for task $k$ at step $t$.
$\tilde{s}_{t}^{k},\tilde{f}_{t}^{k}$ Pseudo success and failure counts from implicit evidence.
Interpolation-based Implicit Evidence
$\bar{p}_{w}^{k},\bar{p}_{s}^{k}$ Pre-computed success rates of a weak and a strong reference model on task $k$.
$\bar{p}_{t}^{\mathrm{ref}},\bar{p}_{w}^{\mathrm{ref}},\bar{p}_{s}^{\mathrm{ref}}$ Avg. success rates of current, weak, and strong models on a reference batch.
$\mu_{t}$ Relative capability coefficient of the current model at step $t$.
$\tilde{\mu}_{t}$ Momentum-updated relative capability coefficient.
$\gamma$ Momentum coefficient for updating $\tilde{\mu}_{t}$.
$\tilde{p}(k,\mathcal{B}_{t})$ Estimated success probability for task $k$ using implicit evidence.
Task Selection (Thompson Sampling)
$p^{*}$ The target success rate for optimal learning (e.g., 0.5).
$\hat{p}_{k}$ Success rate sampled from the posterior $\mathrm{Beta}(\alpha_{t}^{k},\beta_{t}^{k})$
$\hat{u}_{k}$ Utility of a task, computed as $|\hat{p}_{k}-p^{*}|$.
Evaluation Metrics
$\tau$ Target fraction (e.g., 0.5, 0.75) for defining a performance milestone.
$P_{\text{init}},P_{\text{best}}$ Initial and best performance of the baseline method.
$P_{\tau}$ Target performance, $P_{\text{init}}+\tau\cdot(P_{\text{best}}-P_{\text{init}})$.
$\tau_{M}$ Training step (hitting time) for method $M$ to first reach performance $P_{\tau}$.
$\beta$ Budget ratio (e.g., 0.25, 0.5) for evaluating Best-so-far (BSF).
Generalizations & Alternatives (Appendix)
$\mathcal{K}(\cdot,\cdot)$ Kernel function between two tasks.
$\tau$ Temperature parameter for the kernel function.
$\eta$ Natural parameter of an exponential family.
$T(r),A(\eta)$ Sufficient statistic and log-partition function of an exponential family.
$\chi,\nu$ Hyperparameters of the conjugate prior for an exponential family.
Table 5: Symbol Notation
Appendix B Implementational Details
B.1 Training Data
The training data for our experiments is sourced from GURU (cheng2025revisiting), a well-curated, cross-domain RL dataset covering mathematics, code, logic, science, simulation, and tabular tasks. Each subset has been rigorously deduplicated, verified, and filtered. Our training utilizes the math, code, and logic subsets, with the Zebra Puzzle excluded from the logic portion due to its non-binary reward structure.
Mathematics (54.4k)
[…]
Code (18.1k)
The Code subset includes programming challenges from LeetCode (xia2025leetcodedataset), TACO-verified (likaixin2024taco-verified), PrimeIntellect (2025synthetic1), and LiveCodeBench (jain2024livecodebench). For PrimeIntellect and LiveCodeBench, it incorporates pre-filtered versions provided by DeepCoder [^2] https://www.together.ai/blog/deepcoder.
Logic (5.0k)
The Logic subset is composed of several benchmarks designed to test structured reasoning. It includes Ordering Puzzles (relational ordering), Graph Puzzles (implicit graph traversal), the public training splits of ARC-AGI (chollet2024arc) and ARC-AGI-2 (chollet2025arc) (abstract grid transformations), and a 3.4k-sample from BARC (li2024barc) (synthetic ARC-style tasks). For these tasks, predictions are extracted from <answer> tags for reward calculation via exact match.
B.2 Evaluation Benchmarks
Followed by GURU (cheng2025revisiting), the evaluation suite consists of a set of established benchmarks to rigorously assess the model’s performance in a zero-shot setting across the same domains.
Math We evaluate mathematical reasoning on AIME24 (aime2024) and MATH500 (hendrycks2021measuringmathematicalproblemsolving), which cover a wide range of competition-level math problems.
Code Programming capabilities are assessed using HumanEval (chen202humaneval), MBPP (austin2021mbpp), and a subset of LiveCodeBench (jain2024livecodebench). These benchmarks span from basic function generation to complex algorithmic challenges.
Logic Logical and abstract reasoning are measured using Ordering Puzzles for general reasoning, and ARC-AGI (chollet2024arc) for grid-based abstract reasoning.
B.3 Training Details
The RFT setup follows standard practices (cheng2025revisiting; qu2025can; sun2025improving), employing GRPO (grpo2024) as the RL algorithm (detailed in Appendix B.5). All experiments are conducted with 8 NVIDIA A100 (80GB) GPUs. We utilize PyTorch’s FSDP for distributed training, with vLLM (kwon2023efficient) employed to accelerate the response generation during the rollout phase. The actor model is optimized using a learning rate of $1\times 10^{-6}$ and weight decay of 0.1. We apply a constant learning rate warmup for the first 10 steps and use gradient clipping with a maximum norm of 1.0. For the GRPO algorithm, we set the clipping ratio $\epsilon$ to 0.2 and generate 16 rollouts with a temperature of 1.0. The maximum prompt and response lengths are set to 4,096 and 8,192 tokens, respectively.
The training configurations are tailored for two LLMs: Qwen2.5-1.5B-Instruct and Qwen2.5-7B (Yang2024Qwen25TR), across three reasoning domains: Math, Code, and Logic. The primary differences across settings lie in batch sizes and memory optimization strategies to accommodate different model sizes. Specifically, the 7B model experiments utilize a larger training batch size (512 vs. 256) and GRPO mini-batch size (64 vs. 32). To fit the larger model on the same hardware, we enable FSDP’s CPU offloading for all 7B model experiments. Key hyperparameters for all six experimental settings are summarized in Table 6.
Hyperparameter 1.5B-Math 7B-Math 1.5B-Code 7B-Code 1.5B-Logic 7B-Logic
Base Model Qwen2.5-1.5B-Inst. Qwen2.5-7B Qwen2.5-1.5B-Inst. Qwen2.5-7B Qwen2.5-1.5B-Inst. Qwen2.5-7B
Optimizer AdamW AdamW AdamW AdamW AdamW AdamW
Learning Rate $\times 10^{-6} 1×1$$\times 10^{-6}$ $\times 10^{-6} 1×1$$\times 10^{-6}$ $\times 10^{-6} 1×1$$\times 10^{-6}$
Weight Decay 0.1 0.1 0.1 0.1 0.1 0.1
Global Steps 100 100 70 70 100 100
Batch Size (Prompts) 256 512 256 512 256 512
GRPO Mini-batch Size 32 64 32 64 32 64
Rollouts per Prompt ($n$) 16 16 16 16 16 16
GRPO Clip Ratio ($\epsilon$) 0.2 0.2 0.2 0.2 0.2 0.2
Rollout Temperature 1.0 1.0 1.0 1.0 1.0 1.0
Max Prompt Length 4,096 4,096 4,096 4,096 4,096 4,096
Max Response Length 8,192 8,192 8,192 8,192 8,192 8,192
FSDP CPU Offload Disabled Enabled Disabled Enabled Disabled Enabled
Table 6: Key hyperparameters for RFT across different models and domains. “Inst.” is an abbreviation for Instruct.
B.4 Formal Definitions of Metrics
We provide formal definitions of the two key metrics used in our experiments: Time-to-Baseline (TTB) and Best-so-far (BSF).
Time-to-Baseline (TTB).
We define Time-to-Baseline (TTB) as a metric to measure the acceleration of a method relative to the random baseline. Let the random baseline start from performance $P_{\text{init}}$ and reach its best performance $P_{\text{best}}$ within a fixed training window of $K$ steps. For a target fraction $\tau\in\{50\%,75\%,100\%\}$, the target performance is defined as the corresponding interpolation between the initial and best performance:
$P_{\tau}\;=\;P_{\text{init}}+\tau\cdot\big(P_{\text{best}}-P_{\text{init}}\big).$
Denote by $\tau_{M}$ the (possibly interpolated) training step at which method $M$ first achieves performance $P_{\tau}$. Then the TTB of method $M$ is
$\text{TTB}_{M}(\tau)\;=\;\frac{\tau_{M}}{\tau_{\text{baseline}}}.$
By definition, $\text{TTB}_{\text{baseline}}(\tau)=1$. Smaller values of TTB indicate that a method reaches the target improvement faster, and thus achieves greater acceleration relative to the baseline.
Example. Suppose the baseline starts at performance $P_{\text{init}}=0.1$ and reaches $P_{\text{best}}=0.3$ within 100 training steps. The total improvement is therefore $0.2$. For $\tau=50\%$, the target performance is
$P_{50\%}=0.1+0.5\times(0.3-0.1)=0.2.$
If the baseline first reaches $0.2$ at step 40, then $\tau_{\text{baseline}}=40$. If another method $M$ reaches $0.2$ earlier, at step 30, then $\tau_{M}=30$, and thus
$\text{TTB}_{M}(50\%)=\frac{30}{40}=0.75.$
This indicates that method $M$ achieves the same relative improvement (50% of the baseline’s maximum gain) using only 75% of the training steps required by the baseline, reflecting a 25% acceleration.
Best-so-far (BSF).
We define Best-so-far (BSF) to measure the relative performance gain of a method against the random baseline under the same training budget. Let the total training window be $T$ steps. At a budget ratio $\beta\in(0,1]$, corresponding to step $t=\lfloor\beta T\rfloor$, denote by $\mathrm{best}_{M}(t)$ the best performance achieved by method $M$ up to step $t$, and by $\mathrm{best}_{\text{rand}}(t)$ the best performance achieved by the random baseline up to the same step. Then the BSF of method $M$ is defined as
$\mathrm{BSF}_{\beta}(M):=\frac{\mathrm{best}_{M}(t)}{\mathrm{best}_{\text{rand}}(t)}.$
By definition, $\mathrm{BSF}_{\beta}(\text{rand})=1$. Larger values indicate that the method has achieved stronger absolute performance under the same budget, i.e., it delivers better best-so-far outcomes than the baseline, not just relative improvement.
Example. Suppose the total training window is $T=100$ steps, and we consider $\beta=0.5$ ($t=50$). If the random baseline achieves $\mathrm{best}_{\text{rand}}(50)=0.4$, while another method $M$ achieves $\mathrm{best}_{M}(50)=0.6$, then
$\mathrm{BSF}_{50\%}(M)=\frac{0.6}{0.4}=1.5.$
This means that by step 50, method $M$ has achieved a best-so-far performance 1.5 times that of the random baseline.
B.5 Policy Optimization Algorithm: GRPO
The policy is optimized using Group Relative Policy Optimization (GRPO) (grpo2024), a policy gradient method that operates without a learned value function. In GRPO, the advantage for a given response is computed by normalizing its reward against the statistics of a group of candidate responses sampled for the same prompt.
Specifically, for each prompt $x$, a set of $G$ responses $Y=\{y_{1},\dots,y_{G}\}$ is sampled from the policy $\pi_{\theta_{\text{old}}}$. After obtaining the reward $R(y_{k})$ for each response, the group-relative advantage $A(y_{k})$ is defined as:
$A(y_{k})=\frac{R(y_{k})-\mu_{Y}}{\sigma_{Y}}$ (6)
where $\mu_{Y}$ and $\sigma_{Y}$ represent the mean and standard deviation of the rewards $\{R(y_{1}),\dots,R(y_{G})\}$ for the group.
This advantage is then incorporated into a clipped surrogate objective function to update the policy parameters $\theta$:
$\mathcal{L}_{\text{GRPO}}(\theta)=\mathbb{E}_{y\sim\pi_{\theta_{\text{old}}}}\left[\min\left(r_{\theta}(y)A(y),\text{clip}(r_{\theta}(y),1-\epsilon,1+\epsilon)A(y)\right)\right]$ (7)
Here, $r_{\theta}(y)=\frac{\pi_{\theta}(y|x)}{\pi_{\theta_{\text{old}}}(y|x)}$ denotes the probability ratio between the current and old policies. The clip function constrains this ratio within the interval $1-\epsilon,1+\epsilon]$, limiting the magnitude of policy updates during optimization.
B.6 Baseline Details
We detail the configurations of the baselines and ablations used for comparison.
  • Random: Tasks are sampled uniformly at random from the entire training pool. This represents a no-curriculum scenario and serves as the fundamental baseline for measuring performance gains.
  • Offline Baseline: Tasks are pre-sorted once from easy to hard based on the success rates of external models (Qwen2.5-7B-Instruct, with Qwen3-8B for tie-breaking). This baseline represents a static curriculum and is used to benchmark our adaptive method against a fixed task sequence.
  • Setting from MoPPS (qu2025can): This ablation uses $\lambda=0.0,\rho=0.0$ with posterior sampling, relying solely on explicit evidence for task selection. Under this configuration, our framework reduces exactly to the setting studied in qu2025can, enabling a direct evaluation of a purely explicit-evidence-based strategy. Note that qu2025can did not study the code or logic domains, nor did they provide principles for determining $\lambda$ ($1-\lambda$ in BOTS). We therefore reuse their hyperparameter choice for math ($\lambda=1$, i.e., $\lambda=0$ in BOTS) when evaluating other domains. This comparison is fair since no hyperparameter tuning is performed for either MoPPS or BOTS.
  • Proxy for DOTS (sun2025improving): This ablation uses $\lambda=1.0,\rho=1.0$ with posterior sampling disabled, so task selection is driven entirely by implicit evidence from our estimator. It serves as a proxy for the approach in sun2025improving, though our construction of pseudo-counts differs, see Appendix E.2 for more details. This baseline allows us to evaluate the long-term efficacy of a strategy that relies almost exclusively on implicit evidence without direct corrective feedback. We note that sun2025improving additionally explored reusing historical trajectories, which lies beyond the scope of this paper. To ensure clarity, we isolate such tricks from DOTS and focus solely on the online task selection component.
Appendix C More Related Works
Reinforcement Finetuning (RFT).
Reinforcement finetuning (RFT) has become a pivotal paradigm for aligning Large Language Models (LLMs) with human preferences and enhancing reasoning capabilities. Early approaches, commonly known as Reinforcement Learning from Human Feedback (RLHF) (ouyang2022rlhf), train a reward model to approximate human preferences and then optimize the policy using reinforcement learning algorithms such as PPO (2017ppo). Subsequent research introduced direct preference optimization methods such as DPO (2023dpo) and its variants KTO (2024kto), ORPO (2024orpo), and SimPO (2024simpo), which bypass the need to train an explicit reward model. More recently, researchers have observed that when rewards are automatically verifiable—such as in mathematics or code generation—RFT can substantially enhance reasoning performance, leading to the emergence of verifiable-RL frameworks like RLVR, where GRPO (grpo2024) serves as a PPO-style adaptation designed for such settings. Building on these foundations, variants like DAPO (yu2025dapo) and GSPO (2025gspo) further refine the optimization dynamics, while other studies explore entropy regularization (2025primeentropy; 2025entropy28) and reward shaping (2025seedgrpo; 2025trinity). Despite these algorithmic advances, a critical bottleneck remains—the data curriculum, i.e., how tasks are selected and presented during RFT, which crucially determines training efficiency.
The Role of Data Curriculum
The importance of data difficulty in RFT is increasingly recognized, leading to the creation of highly challenging benchmarks (2025bigmath; he2025deepmath; gao2025omni). Beyond dataset creation, data selection and curriculum design has also become a central topic of investigation. For Supervised Finetuning (SFT), a variety of data selection and curriculum strategies have been widely studied based on quality (2024deita; 2024superfilter), diversity (2025daar; 2024instag), or pre-assessed difficulty (2025mindgym; 2025limo). However, such static methods are ill-suited for the dynamic nature of RFT. This has motivated a focus on dynamic curricula that adapt to the model’s evolving capabilities, typically by leveraging the notion of task difficulty (cheng2025revisiting; 2025hardexamples; li2025can).
Task Selection Strategies for RFT
Existing task selection strategies for RFT vary in their approach to leveraging task difficulty. An early line of work, directly inspired by curriculum learning, employs offline curricula that schedule tasks along a fixed easy-to-hard trajectory (parashar2025curriculum; shen2025skywork; zhu2025eduflow; wen2025structured; li2025temple). While simple and intuitive, these methods are non-adaptive and cannot respond to the model’s real-time learning progress. To address this, online selection strategies have emerged. A straightforward approach is sampling-based task filtering, which uses extra rollouts to evaluate and discard tasks that are too easy or too hard, thereby incurring significant computational overhead (yu2025dapo; bae2025online). To avoid this cost, recent works attempt to predict task success rates. Some frame task selection as a non-stationary multi-armed bandit (MAB) problem, but often treat tasks as independent arms, thus overlooking cross-task relationships (chen2025self; qu2025can). Others use a small reference set to predict the difficulty of other tasks via similarity kernels, but this still requires extra rollouts and discards valuable historical information (sun2025improving; wang2025dump).
Positioning Our Work.
The current landscape reveals a clear need for a task selection method that is simultaneously adaptive, computationally efficient, and information-complete. Our framework, BOTS, is designed to fill this gap. It introduces a unified Bayesian framework that is fully online and adaptive. Critically, BOTS is the first to jointly incorporate explicit evidence from direct evaluations and implicit evidence inferred from related tasks, without requiring any additional model rollouts. BOTS provides a principled, low-overhead, and effective solution for dynamic task selection in RFT.
Appendix D Proofs
D.1 Proof for Proposition 3.2
[…]
Proof.
Write the Beta densities (up to normalization) as $\pi_{t}(p)\propto p^{\alpha_{t}-1}(1-p)^{\beta_{t}-1}$ and $\pi_{0}(p)\propto p^{\alpha_{0}-1}(1-p)^{\beta_{0}-1}$. Raising these to powers and multiplying by the (tempered) likelihood terms in equation (4) yields
$\pi_{t+1}(p)\ \propto\ p^{(1-\lambda)(\alpha_{t}-1)+\lambda(\alpha_{0}-1)+(1-\rho)s_{t}+\rho\tilde{s}_{t}}\,(1-p)^{(1-\lambda)(\beta_{t}-1)+\lambda(\beta_{0}-1)+(1-\rho)f_{t}+\rho\tilde{f}_{t}}.$
Collecting exponents shows that $\pi_{t+1}$ is Beta with parameters exactly as stated (add +1+1 back to exponents), completing the proof. ∎
D.2 Proof for Proposition 3.4
[…]
Proof.
We drop the task index for clarity. From Equation (1) and Equation (3),
$n_{t+1}=(1-\lambda)n_{t}+\lambda n_{0}+(1-\rho)(s_{t}+f_{t})+\rho(\tilde{s}_{t}+\tilde{f}_{t})=(1-\lambda)n_{t}+\lambda n_{0}+(1-\rho)n\,\mathbb{I}[\mathcal{E}_{t}]+\rho n,$
where $\mathcal{E}_{t}$ is the event that the task receives a direct evaluation at time $t$ (so $s_{t}+f_{t}=n$ iff $\mathcal{E}_{t}$ holds; otherwise $s_{t}+f_{t}=0$). Unrolling the linear recurrence for $t\geq 0$,
$n_{t}$ $=(1-\lambda)^{t}n_{0}+\sum_{u=0}^{t-1}(1-\lambda)^{u}\!\left(\lambda n_{0}+\rho n+(1-\rho)n\,\mathbb{I}[\mathcal{E}_{t-1-u}]\right)$
$=n_{0}+\frac{\rho n}{\lambda}\!\left(1-(1-\lambda)^{t}\right)+(1-\rho)n\sum_{u=0}^{t-1}(1-\lambda)^{u}\,\mathbb{I}[\mathcal{E}_{t-1-u}],$
where we used $\sum_{u=0}^{t-1}(1-\lambda)^{u}\lambda=1-(1-\lambda)^{t}$. Because $\mathbb{I}[\mathcal{E}_{\cdot}]\in\{0,1\}$,
$0\;\leq\;\sum_{u=0}^{t-1}(1-\lambda)^{u}\,\mathbb{I}[\mathcal{E}_{t-1-u}]\;\leq\;\sum_{u=0}^{t-1}(1-\lambda)^{u}=\frac{1-(1-\lambda)^{t}}{\lambda}.$
Multiplying by $(1-\rho)n$ and adding the common term $n_{0}+\frac{\rho n}{\lambda}(1-(1-\lambda)^{t})$ gives the uniform bounds
$n_{0}+\frac{\rho n}{\lambda}\!\left(1-(1-\lambda)^{t}\right)\;\leq\;n_{t}\;\leq\;n_{0}+\frac{n}{\lambda}\!\left(1-(1-\lambda)^{t}\right).$
Since $(1-\lambda)^{t}\to 0$ exponentially, the $\liminf/\limsup$ statements follow immediately.
Moreover, these bounds are tight: if $\mathbb{I}[\mathcal{E}_{t}]\equiv 0$ (never directly evaluated), then $n_{t}=n_{0}+\frac{\rho n}{\lambda}\bigl(1-(1-\lambda)^{t}\bigr)$ and $\lim_{t\to\infty}n_{t}=n_{0}+\frac{\rho n}{\lambda}$; if $\mathbb{I}[\mathcal{E}_{t}]\equiv 1$ (always directly evaluated), then $n_{t}=n_{0}+\frac{n}{\lambda}\bigl(1-(1-\lambda)^{t}\bigr)$ and $\lim_{t\to\infty}n_{t}=n_{0}+\frac{n}{\lambda}$. ∎
Appendix E Extended Discussion
E.1 Impact of $\lambda,\rho$ on Task Selection
Proposition 3.4 characterizes how the hyperparameters $\lambda$ and $\rho$ jointly determine the effective sample size of each task’s Beta posterior. Intuitively, the effective sample size $n_{t}=\alpha_{t}+\beta_{t}$ controls the confidence of the estimated probability of success, and hence governs the tradeoff between exploration and exploitation in task selection. When a task is repeatedly selected and directly evaluated, its posterior accumulates evidence and $n_{t}$ approaches the upper bound $n_{0}+\tfrac{1}{\lambda}n$. Conversely, when a task is rarely selected and only updated by implicit evidence, $n_{t}$ remains close to the lower bound $n_{0}+\tfrac{\rho}{\lambda}n$.
The role of $\lambda$ is to regulate the overall scale of $n_{t}$. A smaller $\lambda$ accelerates the discounting of historical information, leading to smaller effective sample sizes, higher posterior uncertainty, and thus more exploratory behavior in Thompson Sampling. In contrast, a larger $\lambda$ stabilizes the posterior by retaining more historical counts, biasing the strategy toward exploitation of tasks already known to be near the target probability of success.
The parameter $\rho$ specifically controls the lower bound of $n_{t}$. A smaller $\rho$ increases the uncertainty of tasks without direct evaluations, making it more likely for such tasks to be sampled for exploration even when their estimated probability of success is not close to the target. A larger $\rho$, by assigning more pseudo-counts, suppresses this uncertainty and reduces the chance of revisiting unevaluated tasks, favoring exploitation.
Together, $\lambda$ and $\rho$ provide fine-grained control over the exploration–exploitation balance: $\lambda$ sets the overall confidence level of the posterior across all tasks, while $\rho$ tunes the relative confidence between evaluated and unevaluated tasks. This interaction ensures that the framework adaptively selects tasks of appropriate difficulty as the model evolves, maintaining both robustness and efficiency in online task selection.
E.2 Attention-Based Adaptive Difficulty Estimate
Given a batch of online evaluation results $\mathcal{B}_{t}=\{(\mathcal{T}_{\mathcal{B}_{t}[i]},r_{1:m}^{\mathcal{B}_{t}[i]})\}_{i=1}^{|\mathcal{B}_{t}|}$, we need to estimate $\tilde{s}_{t}^{k}$ and $\tilde{f}_{t}^{k}$ for each task $\mathcal{T}^{k}$. Besides the interplation-based estimate discussed in 3.3, another straightforward idea is to construct a kernel-based estimator:
$\tilde{s}_{t}^{k}=\sum_{i=1}^{|B_{t}|}\frac{\mathcal{K}(\mathcal{T}_{i},\mathcal{T}_{k})}{\sum_{j=1}^{|B_{t}|}\mathcal{K}(\mathcal{T}_{j},\mathcal{T}_{k})}\,s_{t}^{i},\qquad\tilde{f}_{t}^{k}=\sum_{i=1}^{|B_{t}|}\frac{\mathcal{K}(\mathcal{T}_{i},\mathcal{T}_{k})}{\sum_{j=1}^{|B_{t}|}\mathcal{K}(\mathcal{T}_{j},\mathcal{T}_{k})}\,f_{t}^{i}.$ (8)
For example, sun2025improving proposed to use an attention-like kernel structure:
$\mathcal{K}(\mathcal{T},\mathcal{T}^{\prime})=\exp\!\left(\frac{\langle\texttt{Embd}(\mathcal{T}),\texttt{Embd}(\mathcal{T}^{\prime})\rangle}{\tau}\right).$ (9)
However, the limitations of such approaches are two-folds: (1) Due to its convex structure, the kernel estimator does not support extrapolation. Specifically, given a batch of tasks with passing rates bounded within $p_{\text{min}},p_{\text{max}}]$, the estimated passing rate for any task is also restricted to this range. (2) Its effectiveness depends heavily on the quality of the task embeddings, which often require additional training, may be high-dimensional, and thus introduce non-negligible storage overhead.
E.3 Interpolation-Based Implicit Evidence
Extrapolation Capability.
Unlike kernel-based estimators (sun2025improving) (see Appendix E.2 for a brief introduction), which are confined to the convex hull of observed tasks, our interpolation-based estimator naturally supports extrapolation. For example, if the reference pair consists of a strong and a stronger model, the capability coefficient $\mu_{t}(\mathcal{B}_{t})$ defined in Section 3.3 may fall outside $[0,1]$, yielding extrapolated predictions. The clipping step in Equation (5) ensures these estimates remain within the feasible range $[0,1]$.
Choice of Reference Models.
The effectiveness of the implicit evidence mechanism depends critically on the choice of weak and strong reference models, which should ideally bracket the expected capability trajectory of the model under training. If the training model is substantially stronger than the strong reference or weaker than the weak reference, our method must rely on extrapolation, whose accuracy may degrade. In practice, we find that standard publicly available checkpoints—Qwen2.5-7B-Instruct (Yang2024Qwen25TR) and Qwen3-32B (yang2025qwen3)—provide a robust and practical basis, even when training significantly smaller models such as Qwen2.5-1.5B-Instruct (Yang2024Qwen25TR). A systematic ablation on the choice of reference models remains an important direction for future work.
Discussion on the Potential Limitations.
The simplicity of our interpolation-based approach, while a key strength, also entails two potential limitations. (i) Expressive power: The linear interpolation assumes a linear relationship between a model’s global capability and its per-task success rate. While our empirical results suggest this is a powerful approximation (see Appendix F.2 for a validation), the true learning dynamics of LLMs may be more complex. (ii) Distributional shift in capability estimation: The capability coefficient $\mu_{t}$ is estimated on the selected batch $\mathcal{B}_{t}$, which is not a uniform sample from the task pool. This introduces a bias, as the model is likely to perform better on this adaptively chosen batch than on the entire dataset. Consequently, $\mu_{t}$ might be an overestimate of the model’s true global capability. However, we argue this bias is not fatal: the primary role of $\mu_{t}$ is to track the progression of the model’s capability, and even a biased estimate can provide a valuable monotonic signal for this purpose, see Section 4.4 for empirical validation.
E.4 Discussion on Future Works
E.4.1 Generalization to Other Reward Distributions
While our main exposition focuses on binary rewards with a Beta-Bernoulli model, the core principles of BOTS extend naturally to any reward distribution within the exponential family that admits a conjugate prior. This generality makes BOTS a versatile blueprint for online task selection algorithms.
The General Framework.
Let the reward $r$ for a task follow a distribution from a one-parameter exponential family: $f(r\mid\eta)=h(r)\exp(\eta T(r)-A(\eta))$, where $\eta$ is the natural parameter and $T(r)$ is the sufficient statistic. The conjugate prior for $\eta$ takes the form $p(\eta\mid\chi,\nu)\propto\exp(\chi\eta-\nu A(\eta))$, where $(\chi,\nu)$ are hyperparameters. Here, $\chi$ can be seen as a pseudo-sum of sufficient statistics from prior observations, and $\nu$ as a pseudo-count of those observations.
Our generalized Bayesian update rule from Eq. equation (4) can be directly mapped to this setting. The update for the posterior hyperparameters $(\chi_{t},\nu_{t})$ becomes:
$\chi_{t+1}$ $=(1-\lambda)\chi_{t}+\lambda\chi_{0}+(1-\rho)T_{\text{explicit}}+\rho T_{\text{implicit}},$ (10)
$\nu_{t+1}$ $=(1-\lambda)\nu_{t}+\lambda\nu_{0}+(1-\rho)n_{\text{explicit}}+\rho n_{\text{implicit}},$ (11)
where $(\chi_{0},\nu_{0})$ are the base prior’s parameters. For $n_{\text{explicit}}$ direct observations $\{r_{i}\}$, the explicit evidence is $T_{\text{explicit}}=\sum_{i=1}^{n_{\text{explicit}}}T(r_{i})$. For implicit evidence, we assume a pseudo-observation of size $n_{\text{implicit}}$ with an estimated total sufficient statistic $T_{\text{implicit}}$. This structure precisely mirrors our update for the Beta parameters, preserving conjugacy across iterations.
Revisiting the Bernoulli Case.
[…]
Example: Gaussian Rewards.
Consider a continuous reward $r$, like a score from a powerful critic model, modeled as $R\sim\mathcal{N}(\mu,\sigma^{2})$ with known variance $\sigma^{2}$. The conjugate prior for the mean $\mu$ is Gaussian, $\mu\sim\mathcal{N}(\mu_{0},\sigma_{0}^{2})$. In the exponential family form, $\eta=\mu/\sigma^{2}$ and $T(r)=r$. The prior hyperparameters are $\chi_{0}=\mu_{0}/\sigma_{0}^{2}$ and $\nu_{0}=\sigma^{2}/\sigma_{0}^{2}$. The BOTS update would apply directly to $(\chi_{t},\nu_{t})$, where $T_{\text{explicit}}$ is the sum of observed rewards and $T_{\text{implicit}}$ is an estimated sum from the interpolator. The posterior for $\mu$ remains Gaussian, allowing for Thompson sampling by drawing a sample of the mean $\hat{\mu}$ and selecting tasks whose $\hat{\mu}$ is closest to some target score $\mu^{*}$.
Example: Categorical Rewards.
For tasks with $K$ discrete outcomes (e.g., multi-level ratings), the reward is a one-hot vector, and the distribution is categorical. The conjugate prior is the Dirichlet distribution, a multivariate generalization of the Beta. BOTS would maintain a vector of $K$ Dirichlet parameters $(\alpha_{1},\dots,\alpha_{K})$ for each task. The update rules would apply component-wise to each $\alpha_{j}$ based on explicit and implicit counts for that outcome.
This generality significantly broadens the applicability of our framework beyond binary success/failure tasks. It provides a principled and extensible blueprint for difficulty-aware online task selection across a wide spectrum of RFT problems involving diverse reward structures.
E.4.2 Self-Adaptive Update Rules
In the main paper, we set the belief update coefficients $\lambda$ and $\rho$ as fixed hyperparameters. However, our empirical study (Section 4.3 and Section 4.2) shows that different settings benefit different training stages: smaller $\lambda$ accelerates adaptation in early training but may cause instability later, while moderate $\rho$ effectively leverages implicit evidence early on but can reduce accuracy in later stages. A natural extension is to design self-adaptive update rules that automatically adjust $\lambda$ and $\rho$ according to training dynamics—for example, by monitoring posterior uncertainty, validation performance, or the variance of estimated success probabilities. Such adaptive schemes would allow BOTS to dynamically balance exploration and exploitation, potentially improving robustness across diverse tasks and model scales.
E.4.3 Alternative Plug-In for Implicit Evidence
Our interpolation-based estimator provides an extremely lightweight way to generate implicit evidence without additional rollouts, but it is not the only possible choice. More expressive alternatives could be explored, such as kernel-based predictors (sun2025improving), task-embedding regressors, or small auxiliary models trained jointly with the main model. These alternatives may improve predictive accuracy, especially when the reference models poorly bracket the training model’s capability. However, they also introduce a trade-off: richer implicit evidence often requires higher computational and storage costs. Systematically characterizing this trade-off—between accuracy and efficiency in implicit evidence—remains an open research question and a promising direction for future work.
Appendix F Additional Experimental Results
F.1 Computational Overhead
We examine the computational overhead introduced by task selection. The breakdown of wall-clock time across training phases is shown in Figure 4.
Figure 4: Wall-clock time breakdown across training phases for Qwen2.5-1.5B-Instruct (Left) and Qwen2.5-7B (Right) on GURU-Math, trained on 8 A100 GPUs. Runtime is averaged over the first 100 training steps. The cost of task selection—including posterior sampling, index sorting, and distribution parameter updates—is negligible compared to overall training.
As illustrated, the dominant cost arises from generation and model updates, which together account for more than 75% of runtime. By contrast, the overhead of task selection—including posterior sampling, index sorting, and distribution parameter updates—is negligible (0.2% or less). Importantly, unlike generation and model updates, this cost does not increase with model size. Overall, our Bayesian framework and the chosen practical instantiation remain extremely lightweight, adding almost no extra burden to training.
F.2 Interpolation-Based Implicit Evidence: Empirical Results
To empirically validate our interpolation-based implicit evidence estimator, we assess its predictive quality against the evolving empirical success probabilities of the training model. We examine the trajectory under vanilla training, i.e., uniformly sampling tasks for training. At each step, we compare the predictions from our interpolation-based estimator with the ground-truth online task success probabilities. Two metrics are used: (i) Pearson Correlation, measuring the linear relationship between estimated and empirical difficulties, and (ii) ROC AUC, evaluating the ability to distinguish effective tasks (success strictly between 0 and 1) from ineffective tasks (success equal to 0 or 1). Evaluations are conducted throughout training for models of different scales.
Figure 5: Performance of the implicit evidence estimator during training of Qwen2.5-1.5B-Instruct (Left) and Qwen2.5-7B (Right), measured by Pearson Correlation and ROC AUC.
As shown in Figure 5, the estimator exhibits distinct behaviors across model scales. For Qwen2.5-1.5B-Instruct (Left), both Pearson Correlation and ROC AUC remain stable throughout training; the consistently positive correlation and ROC AUC above 0.5 demonstrate that interpolation-based implicit evidence effectively captures task difficulty. Notably, this model is weaker than both reference models (Qwen2.5-7B-Instruct and Qwen3-32B), confirming that our estimator possesses extrapolation capability. For the larger Qwen2.5-7B (Right), both Pearson Correlation and ROC AUC decline in later stages: Pearson Correlation remains positive, but ROC AUC approaches 0.5, indicating that implicit evidence becomes less informative for difficulty prediction as the model matures. This highlights the necessity of incorporating direct evaluations as explicit evidence to maintain accurate estimation.
F.3 Sampling from Posterior
We now investigate the impact of posterior sampling. As discussed in Section 3.4, posterior sampling naturally balances exploration and exploitation in bandit-style problems. Without sampling, tasks with the closest estimated success rates are greedily selected for training, which risks over-exploitation and insufficient exploration.
To examine this effect, we compare our default setting ($\lambda=0.1,\rho=0.1$) with posterior sampling enabled versus disabled. The valid ratio metrics and benchmark performance metrics are reported in Figure 6 and Table 7.
Figure 6: Qwen2.5-1.5B-Instruct on Math. Ratio of sampled training tasks (measured over 16 rollouts) with passing rates: strictly between 0 and 1, strictly greater than 0, and strictly less than 1, along with aggregated performance (MATH500 and AIME24), plotted against training steps.
Benchmark MATH500 AIME24 Math Aggregated Performance
Metric TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$)
Method ($\downarrow$), $\%$ ($\rightarrow$) 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100%
Random 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
$\lambda=0.1,\rho=0.1,$ sample=True 0.89 0.49 0.57 1.13 1.05 1.05 0.51 1.00 1.00 1.25 1.75 1.00 0.89 0.56 0.64 1.12 1.07 1.05
$\lambda=0.1,\rho=0.1,$ sample=False 0.73 0.54 0.54 1.10 1.09 1.03 0.77 0.92 1.25 1.25 1.50 1.00 0.79 0.55 0.81 1.12 1.10 1.07
Table 7: Qwen2.5-1.5B-Instruct on Math. TTB (lower better) and BSF (higher better) evaluated on MATH500, AIME24, and aggregated performance. For TTB, notation ”-” indicates the the target performance is never achieved within the evaluation window. The best and second best results are marked accordingly.
Observations.
When posterior sampling is disabled, the valid ratio exhibits a faster and higher boost in the early phase due to the removal of randomness, but fluctuations appear as training progresses. In contrast, enabling posterior sampling yields a smoother valid ratio trajectory over time. Notably, these differences in valid ratio do not translate into significant differences in benchmark performance: both settings outperform the random baseline and achieve very similar performance levels. Given the improved stability of the valid ratio, we recommend enabling posterior sampling, though this conclusion is less pronounced compared to the effects of $\lambda$ and $\rho$. We leave a larger-scale study for future work to obtain more reliable evidence.
F.4 Dynamics of Selected Tasks
In addition to reporting the Effective Task Ratio (ETR), which reflects the proportion of selected tasks with success probabilities strictly between 0 and 1, we conduct a finer-grained analysis to capture more detailed dynamics. Specifically, we visualize the distribution of success probabilities for selected tasks along the training trajectory, for both Qwen2.5-1.5B-Instruct and Qwen2.5-7B models. This analysis complements ETR by revealing how the quality of selected tasks evolves beyond the binary effective/ineffective distinction.
We use a heatmap where the x-axis represents training steps, the y-axis represents the empirical success rate (discretized from 0/16 to 16/16), and the color intensity indicates the proportion of tasks sampled at that success rate. This allows us to compare the distributional dynamics of BOTS against the random baseline.
Figure 7: Heatmap visualizing the distribution of sampled task success rates over training steps for Random sampling (Top) and BOTS (Bottom) on Qwen2.5-1.5B-Instruct.
The resulting heatmaps in Figure 7 reveal starkly different behaviors. The random baseline (Top) exhibits a largely static distribution, with a persistent, high-density band at the 0/16 success rate, indicating continuous wasted computation on unsolvable tasks. In contrast, BOTS (Bottom) demonstrates a highly dynamic curriculum. The initial concentration of tasks at the 0/16 success rate diminishes rapidly. Concurrently, the sampling density shifts upward, concentrating in the intermediate difficulty range.
This visualization directly illustrates how BOTS actively filters out overly easy or hard tasks and focuses computational resources on the most informative ones, which explains the superior Effective Task Ratio and overall performance gains observed previously.
Figure 8: Heatmap visualizing the distribution of sampled task success rates over training steps for Random sampling (Top) and BOTS (Bottom) on Qwen2.5-7B.
The analysis on the 7B model, shown in Figure 8, reinforces our findings. Consistent with the 1.5B results, BOTS (Bottom) rapidly diminishes sampling of unsolvable tasks (0/16 success rate) and progressively shifts its focus to the intermediate difficulty range. However, an interesting phenomenon emerges in the later training stages due to the 7B model’s stronger capability. For the random baseline (Top), a high-density band appears at the 16/16 success rate, indicating that significant computation is wasted on tasks the model has already mastered. In stark contrast, BOTS effectively avoids this region, maintaining a broad distribution across the intermediate success rates.
This demonstrates BOTS’s advanced adaptivity: it not only filters out tasks that are too hard but also dynamically avoids those that become too easy, thereby maximizing learning efficiency throughout the entire training process.
F.5 Comprehensive Results on Math
Results on Math-1.5B
The results for Qwen2.5-1.5B-Instruct on the GURU-Math dataset are presented in Figure 9, Figure 10, and Table 8. The effectiveness of task selection is first evaluated through the Effective Task Ratio (ETR). BOTS demonstrates a rapid increase in ETR, peaking above 0.8 and sustaining a high level throughout training. In contrast, the random and BOTS-MoPPS baselines show minimal improvement, with ETR remaining below 0.4. While BOTS-DOTS initially performs well, it exhibits a declining ETR in later stages.
This superior task selection efficiency translates into significant gains in downstream performance. The aggregated performance plot shows that BOTS achieves a steeper learning curve and consistently outperforms the baselines. The quantitative results in Table 8 corroborate these findings. Specifically, for the aggregated performance, BOTS achieves a final TTB of 0.64, indicating a 36% reduction in training steps to reach the baseline’s best performance. It also attains a final BSF of 1.05, representing a 5% improvement in peak performance under the same training budget.
Figure 9: Qwen2.5-1.5B-Instruct on Math. Ratio of sampled training tasks with different passing rates.
Figure 10: Qwen2.5-1.5B-Instruct on Math. Performance on downstream code benchmarks (MATH500, AIME24) and their aggregation.
Benchmark MATH500 AIME24 Math Aggregated Performance
Metric TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$)
Method ($\downarrow$), $\%$ ($\rightarrow$) 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100%
Random 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Offline 0.96 0.49 0.72 1.15 1.11 1.02 0.54 - - 1.00 2.00 0.75 0.96 0.51 0.78 1.16 1.12 1.01
BOTS-MoPPS 1.34 1.34 - 0.91 0.90 0.98 0.78 - - 1.25 1.25 0.67 1.78 1.32 - 0.90 0.90 0.99
BOTS-DOTS 0.78 0.51 0.72 1.12 1.10 1.04 0.71 1.12 1.12 0.75 1.50 1.08 0.83 0.75 0.81 1.11 1.09 1.08
BOTS 0.89 0.49 0.57 1.13 1.05 1.05 0.51 1.00 1.00 1.25 1.75 1.00 0.89 0.56 0.64 1.12 1.07 1.05
Table 8: Qwen2.5-1.5B-Instruct on Math. TTB (lower better) and BSF (higher better) evaluated on MATH500, AIME24, and aggregated performance. For TTB, notation ”-” indicates the the target performance is never achieved within the evaluation window. The best and second best results are marked accordingly.
Results on Math-7B
We present the results for the Qwen2.5-7B model on the GURU-Math dataset in Figure 11, Figure 12, and Table 9. Consistent with the 1.5B experiments, our recommended BOTS setting demonstrates robust task selection efficiency. Its ETR rises rapidly to over 0.9 and is maintained at a high level, significantly outperforming the random, offline, and explicit-only baselines. This efficiency translates to tangible performance gains, with Table 9 indicating a final aggregated TTB of 0.63, confirming significant training acceleration.
A notable difference in the 7B setting is the highly competitive performance of BOT-DOTS. Unlike in the 1.5B experiments where its ETR degraded over time, this strategy maintains a high ETR throughout the training process, comparable to that of BOTS. Consequently, its downstream performance is nearly identical to our full BOTS framework, as evidenced by the overlapping performance curves and similar metrics in the corresponding figures and table. This suggests that for a highly capable base model in the math domain, the guidance from implicit evidence alone can be exceptionally effective.
Figure 11: Qwen2.5-7B on Math. Ratio of sampled training tasks with different passing rates over training steps.
Figure 12: Qwen2.5-7B on Math. Performance on downstream math benchmarks (MATH500, AIME24) and their aggregation.
Benchmark MATH500 AIME24 Math
Metric TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$)
Method ($\downarrow$), $\%$ ($\rightarrow$) 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100%
Random 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Offline 0.96 0.70 0.79 1.02 1.01 1.01 0.86 0.90 0.74 0.88 1.00 1.05 0.91 0.73 0.76 0.99 1.00 1.05
BOTS-MoPPS 1.10 1.10 1.35 0.96 1.00 1.00 1.14 0.88 0.64 0.85 1.19 1.10 1.07 1.15 0.70 0.94 1.04 1.04
BOTS-DOTS 0.89 0.72 1.07 1.01 1.00 1.02 0.96 0.76 - 0.89 1.07 0.99 0.83 0.80 0.61 0.98 1.02 1.03
BOTS 0.91 0.74 0.93 1.01 1.02 1.02 0.79 0.63 0.94 0.97 1.11 1.01 0.86 0.77 0.63 0.99 1.04 1.04
Table 9: Qwen2.5-7B on Math. TTB (lower better) and BSF (higher better) evaluated on downstream math benchmarks.
F.6 Comprehensive Results on Code
Results on Code-1.5B
To assess the generalizability of our framework, we replicate the comparative experiments on the GURU-Code dataset. The results, presented in Figure 13, 14 and Table 10, largely affirm the conclusions drawn from the Math domain. Our BOTS again achieves a consistently high ETR, significantly outperforming the random and offline baselines. This demonstrates its robust capability to construct an efficient curriculum across diverse domains. Correspondingly, BOTS delivers strong downstream performance, particularly in the TTB metric, indicating effective training acceleration.
However, the results on Code also reveal distinct characteristics. Most notably, BOTS-DOTS exhibits exceptionally competitive performance. As detailed in Table 10, it matches or even surpasses our recommended BOTS setting on several benchmarks, especially in the BSF metric. This suggests that the implicit evidence from our interpolator is particularly potent for the Code domain. These findings highlight that while the optimal balance of evidence sources may be domain-dependent, BOTS provides a robust and unified solution that remains a top-performing method across both settings.
Figure 13: Qwen2.5-1.5B-Instruct on Code. Ratio of sampled training tasks with different passing rates.
Figure 14: Qwen2.5-1.5B-Instruct on Code. Performance on downstream code benchmarks (HumanEval, MBPP, LiveCodeBench) and their aggregation.
Benchmark HumanEval MBPP LiveCodeBench Aggregated
Metric TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$)
Method ($\downarrow$), $\%$ ($\rightarrow$) 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100%
Random 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Offline 0.69 0.83 1.09 1.12 1.01 1.00 0.69 1.03 - 1.05 0.99 0.99 1.12 1.08 - 0.91 0.88 0.89 0.68 1.03 - 1.09 1.00 0.99
BOTS-MoPPS 0.81 0.77 1.26 1.11 1.02 1.01 0.81 0.81 0.67 1.03 1.02 1.05 0.62 0.62 - 1.05 0.88 0.94 0.78 0.96 1.09 1.09 1.02 1.02
BOTS-DOTS 0.71 0.48 0.54 1.18 1.15 1.01 0.68 0.63 0.89 1.11 1.03 1.03 0.34 1.02 0.85 1.14 0.94 1.22 0.67 0.65 0.79 1.17 1.09 1.02
BOTS 0.56 0.63 - 1.24 1.12 0.99 0.68 0.66 0.67 1.10 1.06 1.05 0.76 1.18 - 1.00 0.76 0.94 0.58 0.90 0.77 1.17 1.08 1.03
Table 10: Qwen2.5-1.5B-Instruct on Code. TTB and BSF evaluated on HumanEval, MBPP, LiveCodeBench, and aggregated performance.
Results on Code-7B
We extended our analysis to the Qwen2.5-7B model to test scalability. The results, presented in Figure 15, 16 and Table 11, primarily demonstrate the consistent robustness of BOTS in optimizing the task selection. As shown in the Valid Ratio plot, our recommended setting achieves the highest proportion of valid tasks, particularly during the early and middle training stages. This consistent efficiency in task selection across different model scales confirms the robustness of our evidence-fusion approach.
On downstream benchmarks, the performance differences between methods are less pronounced, with all strategies converging to a similarly high level of performance. This outcome is expected, as the strong intrinsic capabilities of the 7B base model can diminish the performance differentiation among curriculum strategies on standard benchmarks. Nevertheless, the clear advantage of BOTS in task selection efficiency underscores its value in optimizing computational resources, even when final performance margins are narrow.
Figure 15: Qwen2.5-7B on Code. Ratio of sampled training tasks with different passing rates.
Figure 16: Qwen2.5-7B on Code. Performance on downstream code benchmarks and aggregation.
Benchmark HumanEval MBPP LiveCodeBench Aggregated
Metric TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$)
Method ($\downarrow$), $\%$ ($\rightarrow$) 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100%
Random 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Offline 0.83 1.29 0.92 1.01 0.96 1.00 1.13 1.29 - 0.99 0.97 0.98 1.04 0.96 0.96 0.90 0.98 1.06 0.97 1.24 - 0.99 0.99 0.99
BOTS-MoPPS 0.91 0.54 0.46 1.08 1.05 1.02 0.83 1.21 - 1.03 0.96 0.99 0.97 1.09 - 0.96 0.98 0.96 0.84 0.69 0.84 1.04 1.02 1.00
BOTS-DOTS 0.97 0.80 0.73 1.03 1.03 1.03 0.73 0.68 0.91 1.04 1.00 1.02 0.95 1.32 - 0.96 0.94 0.94 0.86 0.96 0.76 1.02 1.04 1.02
BOTS 0.96 0.55 0.77 1.09 1.03 1.01 0.69 0.87 - 1.03 0.96 1.00 1.18 1.47 - 0.87 0.87 0.98 0.82 0.79 1.06 1.04 1.01 1.00
Table 11: Qwen2.5-7B on Code. TTB and BSF evaluated on downstream code benchmarks.
F.7 Comprehensive Results on Logic
Results on Logic-1.5B
We evaluate the framework’s performance on the GURU-Logic dataset, where our recommended BOTS setting demonstrates competitive performance comparable to its results in the Math and Code domains. The curriculum dynamics in Logic, however, exhibit a unique profile. As shown in Figure 17, the Effective Task Ratio (ETR) for most strategies follows a distinct rise-and-fall pattern.
This task selection behavior is reflected in the downstream performance, detailed in Figure 18 and Table 12. The explicit-only strategy again yields highly competitive results. Against this backdrop, our recommended BOTS setting remains robust, securing strong TTB and BSF scores, particularly in the aggregated performance metrics. These findings underscore that while domain-specific characteristics may favor different evidence sources, BOTS provides a consistently effective and general-purpose solution across all tested environments.
Figure 17: Qwen2.5-1.5B-Instruct on Logic. Ratio of sampled training tasks with different passing rates over training steps.
Figure 18: Qwen2.5-1.5B-Instruct on Logic. Performance on downstream logic benchmarks (Ordering Puzzle, ARCAGI) and their aggregation.
Benchmark Ordering Puzzle ARCAGI Aggregated
Metric TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$)
Method ($\downarrow$), $\%$ ($\rightarrow$) 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100%
Random 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Offline 1.24 1.55 - 0.65 0.78 0.88 1.64 1.16 0.94 0.67 1.11 1.11 1.23 1.36 - 0.67 0.78 0.89
BOTS-MoPPS 0.87 1.02 - 1.12 0.98 0.98 1.23 1.04 0.78 1.00 1.00 1.00 0.87 1.04 - 1.13 0.96 0.96
BOTS-DOTS 0.65 1.31 - 1.27 0.93 0.94 1.00 0.87 - 1.17 0.78 0.89 0.66 1.32 - 1.28 0.92 0.93
BOTS 0.85 0.93 1.03 1.15 1.02 1.02 1.16 0.83 - 1.33 0.89 0.89 0.85 0.94 1.05 1.19 1.01 1.00
Table 12: Qwen2.5-1.5B-Instruct on Logic. TTB and BSF evaluated on downstream logic benchmarks.
Results on Logic-7B
We extend the evaluation to the Qwen2.5-7B model on the GURU-Logic dataset. Consistent with findings in other settings, BOTS demonstrates robust task selection efficiency. As shown in Figure 19, its ETR curve (red line) shows a rapid initial rise, achieving an effective curriculum faster than other methods. On the downstream benchmarks (Figure 20), the performance of all methods converges, which is an expected outcome when finetuning a powerful base model on these tasks.
This setting, however, reveals two interesting phenomena. First, the curriculum dynamics appear highly sensitive to hyperparameter choices. The ETR trajectories for BOTS and BOTS-MoPPS differ markedly, suggesting a complex interplay between evidence sources in this high-capability, specific-domain context. Second, the offline baseline exhibits a distinct, periodic oscillation in its ETR. A plausible hypothesis for this is that the powerful 7B model rapidly masters clusters of easy tasks within the static curriculum, causing ETR to spike, and then encounters clusters of currently insurmountable tasks, causing ETR to plummet. This cycle repeats as the model iterates over the fixed dataset, resulting in the observed sharp fluctuations.
Figure 19: Qwen2.5-7B on Logic. Ratio of sampled training tasks with different passing rates over training steps.
Figure 20: Qwen2.5-7B on Logic. Performance on downstream logic benchmarks and their aggregation.
Benchmark Ordering Puzzle ARCAGI Aggregated
Metric TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$) TTB ($\downarrow$) BSF ($\uparrow$)
Method ($\downarrow$), $\%$ ($\rightarrow$) 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100%
Random 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Offline 1.07 1.21 1.69 0.98 0.97 1.05 1.13 1.81 - 0.71 1.00 0.91 1.06 1.23 0.95 0.96 0.95 1.04
BOTS-MoPPS 0.75 0.89 0.67 1.12 1.00 1.05 0.84 1.47 - 0.71 1.12 0.86 0.75 0.92 0.69 1.07 1.00 1.05
BOTS-DOTS 0.80 0.98 1.61 1.02 0.95 1.03 0.80 0.99 - 1.00 1.00 0.86 0.79 0.91 0.91 1.02 0.95 1.03
BOTS 0.80 0.77 0.89 1.04 1.00 1.00 0.72 1.17 - 0.88 1.12 0.91 0.79 0.78 0.50 1.03 1.01 1.00
Table 13: Qwen2.5-7B on Logic. TTB and BSF evaluated on downstream logic benchmarks.